KyledaGreat:

Find (f∘g)(1) for the following functions. f(−3)=4 and g(1)=−3 (f∘g)(1) =

KyledaGreat:

@vocaloid

Vocaloid:

(f∘g)(x) = f(g(x)) plug x into g(x), plug the end result of g(x) as the x-value of f(x) so in our case (f∘g)(1), we take g(1) = -3 the next step would be to take this output -3, and plug x = -3 into f(x), which they have already said is f(-3) = 4 so (f∘g)(1) = f(g(1)) = f(-3) = 4

KyledaGreat:

1 attachment
KyledaGreat:

i'm thinking the domain would be [4, ∞)

Vocaloid:

yup good for the actual function, you can just add -4x + sqrt(x-4), I don't think you can really simplify beyond that

Vocaloid:

like, you would just write -4x + sqrt(x-4), you can't really combine these terms beyond that

KyledaGreat:

i didn't find no solution

Vocaloid:

@vocaloid wrote:
hm? it's not really asking you to solve anything, just to write the formula for (f+g)(x), which I would personally just write as \[-4x+\sqrt{x-4}\]

KyledaGreat:

correct

KyledaGreat:

same one but different

KyledaGreat:

Consider the following functions. f(x)=−4x and g(x)=√x−4 Find the formula for (f/g)(x) and simplify your answer. Then find the domain for (f/g)(x). Round your answer to two decimal places, if necessary. (f/g)(x) = Domain =

KyledaGreat:

\[\sqrt{x-4} + -4x\] i think this is right

Vocaloid:

(f/g)(x) means divide f(x) by g(x) \[\frac{ -4x }{ \sqrt{x-4} }\] for the domain, remember that 1. the denominator cannot be 0 2. square roots cannot be negative

KyledaGreat:

oh yes ma'am

KyledaGreat:

oh that wasn't right

Vocaloid:

did it tell you which part was wrong? what did you enter for the domain?

KyledaGreat:

[4, ∞)

Vocaloid:

putting sqrt(x-4) in the denominator changes things when sqrt(x-4) was in the numerator, x *could* be exactly 4, because sqrt(4-4) = 0 is valid however, you cannot have x = 4 if sqrt(x-4) is in the denominator, because you cannot divide by 0 so the domain becomes (4, infinity) open bracket around 4, not a closed bracket

KyledaGreat:

alright you was right

KyledaGreat:

Consider the following functions. f={(2,4),(3,−2),(−5,−5)} and g={(−3,1),(−2,5),(3,−2)} Find (f+g)(3) (f+g)(3) =

Vocaloid:

look at the points in the lists f(3) means the value of f at x = 3. this means the x-coordinate is 3. looking at the list of points, we have (3,-2) which means x = 3 and f(3) = -2. so f(3) = -2. repeat this process with the g list. then add the results.

KyledaGreat:

1+4 = 5 -2 + 3 = 1 5 + -2 = 3 3 + -5 = -2 -2 + -5 = -7

Vocaloid:

you only need to find the value of g(3) looking at the list of points in g: g={(−3,1),(−2,5),(3,−2)} (3,-2) is the only point we need, because we only need g(3). (3,-2) means g(3) = -2 so we have f(3) = -2 and g(3) = -2 so (f+g)(3) = -2 - 2 = -4

KyledaGreat:

i can't put the equal sign in there and i see

Vocaloid:

(f+g)(3) = -4, so you would just put -4

KyledaGreat:

Consider the following functions. f={(2,4),(3,−2),(−5,−5)} and g={(−3,1),(−2,5),(3,−2)} (f−g)(3).

KyledaGreat:

(-5, -5)

Vocaloid:

your answer should only be one number we determined that f(3) = -2 and g(x) = -2 so (f-g)(3) is (-2) - (-2) = 0

KyledaGreat:

Consider the following functions. f={(2,4),(3,−2),(−5,−5)} and g={(−3,1),(−2,5),(3,−2)} Find (fg)(3).

Vocaloid:

can you take a screengrab of this one too, I just want to check whether they want the product or the composite

Vocaloid:

is there an open circle between f and g (f⚬g?) or f*g? or just fg?

KyledaGreat:

fg

KyledaGreat:

1 attachment
Vocaloid:

oh in that case you just multiply f(3) * g(3) or just (-2)(-2) = 4

KyledaGreat:

f(3) * g(3) = 9fg that's what i got

KyledaGreat:

but it's probably 4

Vocaloid:

f(3) is not the same as 3*f f(3) means the value of f at x = 3 as we determined before, because (3,-2) is a point on f(x), that means f(3) = -2 likewise, because g has the point (3,-2), that means g(3) = -2 so (fg)(3) = -2 * -2

KyledaGreat:

okay what would i apply because it wasn't right when i put -2 in

Vocaloid:

g(3) is -2 f(3) is -2 which means (fg)(3) = -2 * -2 = 4

KyledaGreat:

Consider the following functions. f={(2,4),(3,−2),(−5,−5)} and g={(−3,1),(−2,5),(3,−2)} Find (f/g)(3).

Vocaloid:

(f/g)(3) is f(3) / g(3) = (-2) / (-2) = 1

KyledaGreat:

Given f(x), find g(x) and h(x) such that f(x)=g(h(x)) and neither g(x) nor h(x) is solely x. f(x)=1/4x+2 g(x) = h(x) =

Vocaloid:

try making h(x) = 1/4x if f(x) = g(h(x)) = 1/4x + 2, what does g(x) need to be?

KyledaGreat:

1/2(2x + 1)

Vocaloid:

hmmm not quite remember, g(h(x)) means take h(x), put it as the x-value into g(x) so if h(x) = (1/4)x, let's try making g(x) = x + 2 if we plug in "(1/4)x" as x into g(x) = x + 2, we get g(h(x)) = (1/4)x + 2, which is the exact f function that we need so one possible answer to this question is h(x) = (1/4)x and g(x) = x + 2

KyledaGreat:

oh okay , i couldn't put (1/4) in

Vocaloid:

\[\frac{ 1 }{ 4 }x\]

KyledaGreat:

it still show it's wrong

Vocaloid:

wait is the denominator 4x + 2?

KyledaGreat:

for gx or hx ?

Vocaloid:

for the original function f(x) can you take a screengrab of this one

KyledaGreat:

oh yes , i'll send u one

KyledaGreat:

1 attachment
Vocaloid:

ok, f(x) = 1 / (4x + 2) , with 4x + 2 in the denominator let's try making h(x) = 4x + 2 in order for g(h(x)) = 1 / (4x + 2), what does g(x) need to be? how do we get the division in the problem?

KyledaGreat:

g(x) needs to be where the number is known to be placed at to be sovled and the division is by hx and gx

Vocaloid:

if f(x) = 1 / (4x + 2) and h(x) = 4x + 2 and f(x) = g(h(x)) = 1 / (4x + 2), that means g(x) = 1/x notice how we can plug in h(x) = 4x + 2 into g(x) to get 1 / (4x + 2)

KyledaGreat:

oh i see g(x) = 1/x

KyledaGreat:

is that true

Vocaloid:

yes. I picked h(x) = 4x + 2 and then I determined that g(x) must be 1/x in order for f(x) = g(h(x)) to be 1 / (4x + 2)

KyledaGreat:

oh okay